# The Algorithmic Regulator

### v2

@@ 1 Introduction
In this paper, we recast the modeling requirement in a setting independent of linearity, probability, or specific signal classes, by using algorithmic information theory (AIT). We model a world $W$ and a regulator $R$ as deterministic causal Turing machines that interact over interface tapes. We denote the world output by $x=O_{W}$ (over some temporal horizon of length $N$). Our main technical claim is that regulation in the algorithmic sense, i.e., simplicity, forces algorithmic dependence between $W$ and $R$.
@@ 4 Discussion
Our results are single-episode and distribution-free: they make statements about an individual realized readout $x$ and about the pair $(W,R)$ as concrete programs, without positing a stochastic source. Classical (Shannon) information theory quantifies expected code lengths and mutual information with respect to a specified probability law; entropy $H(X)$ and mutual information $I(X;Y)$ are undefined without a distribution, and asymptotic statements (AEP/typical sets) further require ergodicity/mixing assumptions (Shannon 1948; Cover–Thomas). In our setting, there is no given probabilistic model over worlds, regulators, or outputs—indeed, the point is to infer model content from a single realized $x$.
@@
Finally, while one can approximate $K(\cdot)$ with MDL/codelengths in practice, MDL’s justification itself rests on the AIT view that shorter descriptions are better and on the coding‑theorem linkage between description length and (universal) probability (Grünwald 2007). In short: AIT provides the universal prior ($m$), object‑level complexities ($K$), and mutual algorithmic information ($M$) needed to turn the informal slogan “see a simple string, suspect a simple generator” into posterior and contrastive theorems—none of which can be stated in Shannon’s framework without ad hoc model classes and priors.
@@
Our statements are thus complementary and distinct: we work in a distribution‑free, program‑level setting and make no linearity or smoothness assumptions. We do not assert the existence of a dynamical replica inside $R$. Instead, we show that sustained contrastive compressibility ($\Delta>0$) tilts the universal posterior toward pairs $(W,R)$ with larger mutual algorithmic information M(W:R)M(W{:}R), i.e., $R$ carries algorithmic structure about $W$. Thus, “the regulator contains a model” is made precise as M(W:R)>0M(W{:}R)>0 (information‑theoretic dependence), not as an embedded exosystem. The IMP supplies structural necessity for perfect regulation within specified signal classes; our AIT results supply information‑theoretic necessity for observed compressibility advantages, beyond linearity or probabilistic assumptions [37].
@@
Exogenous generator Exosystem $E$ = autonomous generator of reference/disturbances; unaffected by $C$. Fold into the World $W$; no architectural split required (you may still conceptually identify this subpart). Reference: setpoint schedule $r(t)$ (often clock‑driven). Disturbances: outdoor temperature, solar load, occupancy heat gains.
Plant Plant $P$ = room thermal dynamics + heater. Also inside World $W$. First‑order building thermal model, heater actuation, heat losses, sensor dynamics.
@@
Measured output $y$. Part of $O_{W}$ (chosen readout). Indoor temperature (or a weighted error signal).
Error / objective e=rye=r-y; IMP concerns asymptotic $e\!\to\!0$ for a signal class. Score regulation by compressibility of a task readout (e.g., $e_{1:T}$) vs. an OFF baseline; gap $\Delta=K(y_{\text{off}})-K(y_{\text{on}})$. Good thermostat $\Rightarrow$ the error stays near a regular pattern (within deadband) $\Rightarrow$ shorter code length than “heater OFF”.
Table 1: Mapping the IMP triple $(E,C,P)$ and the AIT $(W,R)$ formulation to a thermostat. IMP sources: [13, 14, 37]; AIT view: this work.
@@
The theorems are stated in terms of prefix Kolmogorov complexity, which is not computable. In practice, one fixes a reference prefix code $C$ and estimates
@@
Some standard choices are Lempel-Ziv compressors (LZ77/LZ78/LZW). LZ‑type compressors are universal in a weak sense for stationary ergodic sources and are widely available. Implementations (gzip, lz4, etc.) are practical proxies for $L_{C}(\cdot)$ (Ziv–Lempel 1977; Ziv–Lempel 1978). Recommendations. (i) Fix compressor, window size, and container overhead once; report the raw codelength in bits (not just ratios). (ii) Use identical pre‑processing and record layout across ON/OFF runs. (iii) For short $N$, prefer LZ78‑style (dictionary‑based) or arithmetic‑coded LZ with small headers to reduce fixed overhead. (iv) When comparing across tasks, control for compressor statefulness (reset between episodes).
If both ON and OFF strings are available and one wants a scale‑free sanity check of contrast, compute
@@
where $C(\cdot)$ is the chosen code length and $xy$ is concatenation (Cilibrasi–Vitányi 2005). NCD is heuristic but can reveal whether $x$ is “closer” to trivial baselines than $y$.
The Block Decomposition Method (BDM) estimates $K$ by tiling a string (or array) into small blocks whose complexities are looked up from Coding‑Theorem‑Method (CTM) tables (exhaustive output frequency statistics of small machines), plus a logarithmic penalty for multiplicities:
$\widehat{K}_{\mathrm{BDM}}(x)\ \approx\ \sum_{i}K_{\mathrm{CTM}}(b_{i})+\log m_{i},$
where $b_{i}$ are distinct blocks and $m_{i}$ their multiplicities (see [34, 44]). This is sensitive to small‑scale algorithmic regularities beyond LZ’s parse statistics; it works on 1D/2D data (but depends on the chosen CTM table — size and machine model — and it suffers from boundary/tiling effects and additive constants that can be sizable for short $N$).
@@ 5 Conclusion
The result supplies: (i) a distribution‑free, single‑episode diagnostic for “does the controller contain a model?”, (ii) a complement to the IMP (which requires embedding a copy of the signal generator under structured assumptions), and (iii) a simple experimental recipe—fix a lossless compressor, quantize the readout, compute two code lengths (ON vs. OFF), and use their difference $\Delta$ as evidence of model content in the controller.
@@ Acknowledgments
The author thanks Francesca Castaldo for reviewing the manuscript.
@@ A.1 Setting and core definitions
By the Coding Theorem there exist machine–dependent $c_{1},c_{2}>0$ with $c_{1}2^{-K(x)}\leq m(x)\leq c_{2}2^{-K(x)}$ ( [35]; [47]; Vitányi 2013; Li–Vitányi (book)).

### v3

@@ 1 Introduction
In this paper, we recast the modeling requirement in a setting independent of linearity, probability, exact regulation or specific signal classes, by using algorithmic information theory (AIT). We model a world $W$ and a regulator $R$ as deterministic causal Turing machines that interact over interface tapes. We denote the world output by $x=O_{W}$ (over some temporal horizon of length $N$). Our main technical claim is that regulation in the algorithmic sense, i.e., simplicity, forces algorithmic dependence between $W$ and $R$.
@@ 4 Discussion
Our results are single-episode and distribution-free: they make statements about an individual realized readout $x$ and about the pair $(W,R)$ as concrete programs, without positing a stochastic source. Classical (Shannon) information theory quantifies expected code lengths and mutual information with respect to a specified probability law; entropy $H(X)$ and mutual information $I(X;Y)$ are undefined without a distribution, and asymptotic statements (AEP/typical sets) further require ergodicity/mixing assumptions [13]. In our setting, there is no given probabilistic model over worlds, regulators, or outputs—indeed, the point is to infer model content from a single realized $x$.
@@
Finally, while one can approximate $K(\cdot)$ with MDL/codelengths in practice, MDL’s justification itself rests on the AIT view that shorter descriptions are better and on the coding‑theorem linkage between description length and (universal) probability [22]. In short: AIT provides the universal prior ($m$), object‑level complexities ($K$), and mutual algorithmic information ($M$) needed to turn the informal slogan “see a simple string, suspect a simple generator” into posterior and contrastive theorems—none of which can be stated in Shannon’s framework without ad hoc model classes and priors.
@@
The home thermostat.
As an example, consider a home thermostat as regulator/controller. Let $P$ be the living room + heater dynamics (thermal capacitance, heat loss, delays) and $E$ the exogenous processes (setpoint schedule, outdoor weather/solar, occupancy). The Internal Model Principle (IMP) states that exact output regulation for a specified signal class is possible only if the controller embeds a copy of the exosystem $E$ that generates those signals (e.g., an integrator for steps, an oscillator for a fixed sinusoid); plant knowledge is used for stabilization/shaping, but the IMP necessity targets $E$ itself [19, 9]. In our AIT view, a regulator $R$ is “good” when it makes the realized readout more compressible than a null baseline; a sustained compressibility gap implies that $R$ shares computable structure with the whole world W=(P,E)W{=}(P,E):
$M(W{:}R)\;=\;M\big((P,E):R\big)\;=\;M(P{:}R)\;+\;M(E{:}R\mid P)\;\pm O(\log).$
A simple on/off thermostat with a deadband tuned to the room time constant typically yields a bounded limit cycle (not zero steady-state error); under IMP it lacks the needed internal model of constants (no embedded integrator), hence it does not achieve exact regulation of the “constant” class [19, 4]. Nevertheless, in the AIT sense it still qualifies as a regulator: its policy encodes a very compressed model spanning $P$ (heating raises $T$, room inertia) and weak regularities in $E$ (quasi-constant setpoint, slowly varying weather), giving M(W:R)>0M(W{:}R)>0 [12]. PI/PID or predictive thermostats remedy the IMP shortfall by embedding the appropriate internal model (and often explicit models of $P$ and aspects of $E$) [4].
The AIT regulator framework (as well as the original GRT) is therefore more general than IMP: the regulator must carry a model of the world $W=E\cup P$, where $P$ is the plant (house/HVAC thermodynamics) and $E$ the exogenous processes (setpoint schedule, weather/solar/occupancy), and IMP is recovered as a special case when the performance target is exact output regulation over a specified signal class. Under IMP, a controller qualifies for exact regulation only if it embeds a dynamical copy of the exosystem that generates the reference/disturbances (e.g., an integrator for steps, oscillators for sinusoids)—no model of $P$ is required beyond stabilizability/detectability [19]; nonlinear output-regulation extends this under additional immersion/detectability and regulator-equation solvability assumptions [9].
@@
Exogenous generator Exosystem $E$: autonomous generator of references/disturbances (no feedback from $C$); exact regulation is defined w.r.t. a signal class $\mathcal{U}$. Fold into the World $W$; no architectural split is required (but may still conceptually identify this subpart). Reference $r(t)$: setpoint schedule (often clock‑driven). Disturbances: outdoor temperature, solar load, occupancy heat gains.
Plant Plant $P$: room thermal dynamics + actuator/sensor; used for stabilization/shaping. Also inside World $W$. RRCC (thermal) model, heater actuation, heat losses, sensor dynamics/delay.
@@
Measured output $y$. World readout $x$ extracted from the transcript (often x=yx{=}y or the error string $e_{1:T}$). Indoor temperature $T_{\mathrm{in}}$ (or a weighted error signal).
Error / objective e=rye=r-y; IMP concerns asymptotic $e\!\to\!0$ for all $r,d$ in the class $\mathcal{U}$ (internal model must match $E$). Score regulation by compressibility of the chosen readout $x$ with $R$ ON vs. an OFF baseline ($R{=}\varnothing$). Define the gap $\Delta\;=\;K(x_{\text{off}})\;-\;K(x_{\text{on}})$ (practically, use a fixed MDL code $L_{C}$ in place of $K$). Good thermostat $\Rightarrow$ $x_{\text{on}}$ (e.g., temperature or error) stays near a regular deadband pattern $\Rightarrow$ shorter code than the null/open‑loop case (heater OFF or fixed duty).
Table 1: Mapping the IMP triple $(E,C,P)$ and the AIT $(W,R)$ view to a simple thermostat. IMP emphasizes an internal model of the exosystem $E$ for exact regulation over a signal class; AIT treats W=(P,E)W{=}(P,E) jointly and assesses regulation by a compressibility advantage $\Delta$.
Our statements are thus complementary and distinct: in AIT, we work in a distribution‑free, program‑level setting and make no linearity or smoothness assumptions. We remain agnostic about what the regulator needs to model and do not demand exact regulation. We do not assert the existence of a dynamical replica inside $R$. Instead, we show that sustained contrastive compressibility ($\Delta>0$) tilts the universal posterior toward pairs $(W,R)$ with larger mutual algorithmic information M(W:R)M(W{:}R), i.e., $R$ carries algorithmic structure about $W$. Thus, “the regulator contains a model” is made precise as M(W:R)>0M(W{:}R)>0 (information‑theoretic dependence), not as an embedded exosystem. The IMP supplies structural necessity for perfect regulation within specified signal classes; our AIT results supply information‑theoretic necessity for observed compressibility advantages, beyond linearity or probabilistic assumptions [51].
@@
Our theorems are stated in terms of prefix Kolmogorov complexity, which is not computable. In practice, one can fix a reference prefix code $C$ and estimate upper bounds,
@@
Some standard choices for providing upper bounds to Kolmogorov complexity are Lempel-Ziv compressors (LZ77/LZ78/LZW). LZ‑type compressors are universal in a weak sense for stationary ergodic sources and are widely available. Implementations (gzip, lz4, etc.) are practical proxies for $L_{C}(\cdot)$ [63, 42]. If both ON and OFF strings are available and a scale‑free sanity check of contrast is needed, we can compute
@@
where $C(\cdot)$ is the chosen code length and $xy$ is concatenation [32, 11]. NCD is heuristic but can reveal whether $x$ is “closer” to trivial baselines than $y$.
The Block Decomposition Method (BDM) estimates $K$ by tiling a string (or array) into small blocks whose complexities are looked up from Coding‑Theorem‑Method (CTM) tables (exhaustive output frequency statistics of small machines), plus a logarithmic penalty for multiplicities, $\widehat{K}_{\mathrm{BDM}}(x)\ \approx\ \sum_{i}K_{\mathrm{CTM}}(b_{i})+\log m_{i},$ where $b_{i}$ are distinct blocks and $m_{i}$ their multiplicities (see [48, 61]). This is sensitive to small‑scale algorithmic regularities beyond LZ’s parse statistics; it works on 1D/2D data (but depends on the chosen CTM table — size and machine model — and it suffers from boundary/tiling effects and additive constants that can be sizable for short $N$).
Finally, alternatives include learned compressors based on neural networks. Autoencoder/variational–autoencoder codecs optimize a rate–distortion (thus MDL) objective, with an explicit codelength view via ELBO and practical lossless coding through bits‑back [22, 24, 31, 54, 25]. In images and video, end‑to‑end trained autoencoders, hyperpriors, and autoregressive priors are now standard [6, 37, 23]. Transformer‑based compressors are competitive and rapidly improving—both for images via transformer or hybrid Transformer–CNN codecs [36, 35] and for general lossless compression using language‑model predictors and specialized transformer compressors [15, 55]; see also cross‑modal results reported with large models [34] and neural codecs for audio [60]. From an MDL perspective, these models implement universal codes whose lengths upper‑bound the negative log‑likelihood under the learned generative model.
@@ 5 Conclusion
The result supplies: (i) a distribution‑free, single‑episode diagnostic for “does the controller contain a model?”, (ii) a complement to the IMP (which requires embedding a copy of the signal generator under more restrictive and structured assumptions), and (iii) a simple experimental recipe—fix a lossless compressor, quantize the readout, compute two code lengths (ON vs. OFF), and use their difference $\Delta$ as evidence of model content in the controller.
@@ Acknowledgments
The author thanks Francesca Castaldo for discussions and reviewing the manuscript. The author thanks David Wolpert (SFI) for highilighting open questions regarding the classical regulator theorem.
@@ A.1 Setting and core definitions
By the Coding Theorem there exist machine–dependent $c_{1},c_{2}>0$ with $c_{1}2^{-K(x)}\leq m(x)\leq c_{2}2^{-K(x)}$ [49, 64, 33, 58].